Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $3.40 per unit. Enough capacity exists in the company’s plant to produce 30,500 units of the toy each month. Variable expenses to manufacture and sell one unit would be $2.14, and fixed expenses associated with the toy would total $57,145 per month.

The company's Marketing Department predicts that demand for the new toy will exceed the 30,500 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $2,857 per month. Variable expenses in the rented facility would total $2.38 per unit, due to somewhat less efficient operations than in the main plant.

Required:

1. What is the monthly break-even point for the new toy in unit sales and dollar sales.

2. How many units must be sold each month to attain a target profit of $13,056 per month?

3. If the sales manager receives a bonus of 25 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 25% return on the monthly investment in fixed expenses
Answer #1

Page No 0 Answer © Breaks even point in mat sales and dollar sales Contribution margin per unit = $3:40 – $2.14 = $p. 26 perPage No faced expenses = $18715 + $2854 . - $ 21542 21572 No. of units = 70a R = 21149 units Break even point in vnit sales =Page: No ® 21572 +13056 102 34628 1.02 = 33949 units units to be sold = 30500 + 33949 = 64449 units comits to be sold = 64449Page: No 9 Contribution mwigin per unit - $3.40 - $ 2.63 = $0.97 No. of units a $15001 0.77 19481.818 x 19482 units units to